### v1

@@ 5.4 LLMs/VLMs-based Methods
Wu et al. [328] present a vision language task called Referring Multi-Object Referring (RMOT), which enables queries in natural language to refer to and track multiple objects over time in video sequences. RMOT supports flexible object cardinality and dynamic temporal variations in object status. Authors develop a benchmark, Refer-KITTI, based on the KITTI dataset, providing high-quality annotations with low labeling cost and supporting multi-object, temporally-aware queries. Additionally, their proposed TransRMOT uses an early-fusion cross-modal encoder and a query-decoupled decoder for simultaneous detection and tracking. TransRMOT performs better than CNN-based and Transformer-based baselines on the Refer-KITTI benchmark, highlighting the RMOT formulation’s effectiveness and the proposed architecture for real-world, language-guided multi-object tracking. The authors in [319] propose a large-scale benchmark to combine natural language prompts into 3D perception for autonomous driving. NuPrompt is based on the nuScenes dataset, which contains more than 40,000 finely annotated language-object pairs, with each prompt referring to multiple object trajectories across 3D, multiview, and temporal frames. For simplicity, the authors define a new task for predicting object motion based on descriptive prompts and propose a baseline model, PromptTrack, incorporating a novel prompt reasoning module into a transformer-based tracking architecture. The results show that PromptTrack outperforms several strong baselines and validates the effectiveness of prompt-guided object tracking and forecasting. Their method is the foundation for research at the intersection of vision-language modeling and AV systems.
@@
Chen et al. [329] propose a framework, Contrastive Language Image Pretraining (CLIP) models, for 3D scene understanding tasks, particularly for semantic segmentation of point clouds. They address the challenges of limited 3D annotations and domain mismatch between 2D and 3D data. Also, the authors propose a Semantic-driven Cross-modal Contrastive Learning strategy combining semantic consistency and spatial-temporal consistency. This method enables effective annotation-free segmentation, achieving state-of-the-art results on datasets like nuScenes and ScanNet, and also shows better performance in few-shot learning scenarios. This work is among the first to bridge 2D vision language pre-training and 3D scene understanding through cross-modal contrastive learning. Taparia et al. [317] present a fusion method that uses VLMs to improve object detection under various environmental conditions. Some fusion methods often apply static fusion rules and have challenges in unseen scenarios. At the same time, VLC Fusion addresses this limitation by conditioning the fusion process on environmental context extracted using large VLMs such as GPT-4o. The authors introduce an automated pipeline to extract and query application-specific ecological cues, which are then used to modulate multimodal feature fusion via Feature-wise Linear Modulation (FiLM). Experiments on Waymo Open (RGB + LiDAR) and ATR (visible + infrared) datasets show that VLC Fusion outperforms other baselines.

### v2

@@ 5.4 LLMs/VLMs-based Methods
Wu et al. [wu2023referring] present a vision language task called Referring Multi-Object Referring (RMOT), which enables queries in natural language to refer to and track multiple objects over time in video sequences. RMOT supports flexible object cardinality and dynamic temporal variations in object status. Authors develop a benchmark, Refer-KITTI, based on the KITTI dataset, providing high-quality annotations with low labeling cost and supporting multi-object, temporally-aware queries. Additionally, their proposed TransRMOT uses an early-fusion cross-modal encoder and a query-decoupled decoder for simultaneous detection and tracking. TransRMOT performs better than CNN-based and Transformer-based baselines on the Refer-KITTI benchmark, highlighting the RMOT formulation’s effectiveness and the proposed architecture for real-world, language-guided multi-object tracking. The authors in [wu2025language] propose a large-scale benchmark to combine natural language prompts into 3D perception for autonomous driving. NuPrompt is based on the nuScenes dataset, which contains more than 40,000 finely annotated language-object pairs, with each prompt referring to multiple object trajectories across 3D, multiview, and temporal frames. For simplicity, the authors define a new task for predicting object motion based on descriptive prompts and propose a baseline model, PromptTrack, incorporating a novel prompt reasoning module into a transformer-based tracking architecture. The results show that PromptTrack outperforms several strong baselines and validates the effectiveness of prompt-guided object tracking and forecasting. Their method is the foundation for research at the intersection of vision-language modeling and AV systems.
@@
Chen et al. [chen2023clip2scene] propose a framework, Contrastive Language Image Pretraining (CLIP) models, for 3D scene understanding tasks, particularly for semantic segmentation of point clouds. They address the challenges of limited 3D annotations and domain mismatch between 2D and 3D data. Also, the authors propose a Semantic-driven Cross-modal Contrastive Learning strategy combining semantic consistency and spatial-temporal consistency. This method enables effective annotation-free segmentation, achieving state-of-the-art results on datasets like nuScenes and ScanNet, and also shows better performance in few-shot learning scenarios. This work is among the first to bridge 2D vision language pre-training and 3D scene understanding through cross-modal contrastive learning. Taparia et al. [taparia2025vlc] present a fusion method that uses VLMs to improve object detection under various environmental conditions. Some fusion methods often apply static fusion rules and have challenges in unseen scenarios. At the same time, VLC Fusion addresses this limitation by conditioning the fusion process on environmental context extracted using large VLMs such as GPT-4o. The authors introduce an automated pipeline to extract and query application-specific ecological cues, which are then used to modulate multimodal feature fusion via Feature-wise Linear Modulation (FiLM). Experiments on Waymo Open (RGB + LiDAR) and ATR (visible + infrared) datasets show that VLC Fusion outperforms other baselines.